# Diffusion Autoencoders for Few-shot Image Generation in Hyperbolic Space

### v1

@@ 10 Out-of-distribution Few-shot Image Generation
In Sec 4.2 of the main paper, we mentioned we fine-tuned the model trained with VGGFaces using the FFHQ dataset. The model shows exceptional out-of-distribution generalization ability on the FFHQ dataset. To further verify the OOD generalization ability of HypDAE, we select two images for Animal Faces [45], Flowers [51], and NABirds [67] datasets with three styles from DomainNet [54] including “ painting”, “sketch”, “quick draw”, and “clipart” styles where are model never seen during the training stage. The OOD style transfer can be done by slightly increasing the encoding strength of the stochastic encoder to capture more style information of the given new images. The results in Fig. 18 Fig. 19 Fig. 20 Fig. 21 show that our proposed method has exceptional OOD generalization ability even for new domains with a big gap with the original domain. Although our model still generates images with some real detail for the style “clipart”, the performance in other styles is satisfying. Such an OOD generalization ability is significantly better than any of the previous works before.
@@ 11 Hierarchical Image Generation
In this section, we provide additional examples of images generated by HypDAE at varying radii in the Poincaré disk. As illustrated in Fig. 22Fig. 23, and Fig. 24, high-level, category-relevant attributes remain unchanged when the radius is large, allowing for the generation of diverse images within the same category. Conversely, as the hyperbolic radius $r_{\mathbb{D}}$ decreases, the generated images become more abstract and semantically diverse. Moving closer to the center of the Poincaré disk results in the gradual loss of fine-grained details and changes to higher-level attributes.
@@ 12 Comparison with State-of-the-art Few-shot Image Generation Method
We compare images generated by state-of-the-art methods, including WaveGAN [71], HAE [41], and our proposed method, across four datasets. As shown in Fig. 25, WaveGAN produces high-fidelity images, but the diversity is limited (i.e., blending features from two input images without significant variation). HAE improves diversity but suffers from low fidelity and quality, with missing details and changes in category or identity compared to the original images. In contrast, HypDAE achieves an excellent balance between maintaining identity and enhancing diversity while delivering significantly higher image quality than other methods. These results highlight the potential of HypDAE for broader applications in future downstream tasks.
@@ 13 User Study
The user-study interface is shown in Fig. 28.
@@ 14 Additional Examples Generated by HypDAE
Finally, we provide more examples generated by HypDAE in Fig. 26 and Fig. 27 for four datasets. The results show that our method achieves a balance between the quality and diversity of the generated images which significantly outperforms previous methods.

### v2

@@ 11 Out-of-distribution Few-shot Image Generation
In Sec 4.2 of the main paper, we mentioned we fine-tuned the model trained with VGGFaces using the FFHQ dataset. The model shows exceptional out-of-distribution generalization ability on the FFHQ dataset. To further verify the OOD generalization ability of HypDAE, we select two images for Animal Faces [39], Flowers [43], and NABirds [58] datasets with three styles from DomainNet [46] including “ painting”, “sketch”, “quick draw”, and “clipart” styles where are model never seen during the training stage. The OOD style transfer can be done by slightly increasing the encoding strength of the stochastic encoder to capture more style information of the given new images. The results in Fig.˜20 Fig.˜21 Fig.˜22 Fig.˜23 show that our proposed method has exceptional OOD generalization ability even for new domains with a big gap from the original domain. Although our model still generates images with some real detail for the style “clipart”, the performance in other styles is satisfying. Such an OOD generalization ability is significantly better than any of the previous works.
@@ 12 Hierarchical Image Generation
In this section, we provide additional examples of images generated by HypDAE at varying radii in the Poincaré disk. As illustrated in Fig.˜24Fig.˜25, and Fig.˜26, high-level, category-relevant attributes remain unchanged when the radius is large, allowing for the generation of diverse images within the same category. Conversely, as the hyperbolic radius $r_{\mathbb{D}}$ decreases, the generated images become more abstract and semantically diverse. Moving closer to the center of the Poincaré disk results in the gradual loss of fine-grained details and changes to higher-level attributes.
@@ 13 Comparison with State-of-the-art Few-shot Image Generation Method
We compare images generated by state-of-the-art methods, including WaveGAN [61], HAE [37], and our proposed method, across four datasets. As shown in Fig.˜27, WaveGAN produces high-fidelity images, but the diversity is limited (i.e., blending features from two input images without significant variation). HAE improves diversity but suffers from low fidelity and quality, with missing details and changes in category or identity compared to the original images. In contrast, HypDAE achieves an excellent balance between maintaining identity and enhancing diversity while delivering significantly higher image quality than other methods. These results highlight the potential of HypDAE for broader applications in future downstream tasks.
@@ 14 User Study
The user-study interface is shown in Fig.˜30.
@@ 15 Additional Examples Generated by HypDAE
Finally, we provide more examples generated by HypDAE in Fig.˜28 and Fig.˜29 for four datasets. The results show that our method achieves a balance between the quality and diversity of the generated images which significantly outperforms previous methods.